OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4.

OpenStudy (anonymous):

i think its y = 2x - 3

OpenStudy (anonymous):

Ⓐ Loujoelou

OpenStudy (aroub):

y=mx+b; (1,-5) , m=2 -5=2(1)+b -5=2+b -7=b => y=2x-7 I think your mistake was in this step => "-5=2+b" you have to subtract 2 to both sides, -5-2=7

OpenStudy (anonymous):

y = -2x + 7

OpenStudy (anonymous):

@aroub

OpenStudy (anonymous):

@aroub

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@myininaya